# SMReferences for Appendix

### v3

@@ A.5 Proof of Theorem 5
We begin with the following lemma on the negligible effect of trimming in $\hat{\pi}$ when $s_{\pi}\ll\sqrt{n}/\log p$:
@@
Under the conditions of Theorem 5, we have that with probability converging to $1$, for all $i$,
$\hat{\pi}_{i}\equiv\phi(X_{i}^{\top}\hat{\gamma})\quad\&\quad\hat{\pi}_{\mathrm{aux},i}\equiv\phi(X_{\mathrm{aux},i}^{\top}\hat{\gamma})$
@@
where for the last inequality we use Condition 8 and that $s_{\pi}=o(\sqrt{n}/\log p)$. Thus the statement of this lemma is proved. ∎
Armed with the above lemma, we can see that in the large sample limit, $\hat{\tau}_{\textrm{SCal},r}$ is numerically equivalent to the estimator by redefining $\hat{\pi}\coloneqq\phi(X^{\top}\hat{\gamma})$, which in turn makes $\hat{\tau}_{\textrm{SCal},r}$ numerically equivalent to a variant of the DIPW estimator [Wang and Shah, 2020] in asymptotic sense. This allows us to obtain the following intermediate result on the asymptotic property of $\hat{\tau}_{\textrm{SCal},r}$:
Theorem 10.
[Wang and Shah, 2020, Theorem 2] Under the same conditions of Theorem 5, we have that $\sqrt{n}(\hat{\tau}_{\textrm{SCal},r}-\bar{\tau}^{\ast})$ has the same asymptotic representation of $\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast})$ as stated in Theorem 5.
Speculating the proof in Theorem 2 of Wang and Shah [2020], it boils down to controlling the difference between $\hat{\tau}_{\textrm{DCal}}$ and $\hat{\tau}_{\textrm{SCal},r}$:
$\hat{\tau}_{\textrm{DCal}}-\hat{\tau}_{\textrm{SCal},r}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\pi}_{i}}-\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\hat{\pi}_{i}}\right)$
$=\underset{\mathrm{I}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}\tilde{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})}}-\underset{\mathrm{II}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}(\hat{\pi}_{i}-\tilde{\pi}_{i})}}-\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-\frac{T_{i}}{\pi_{i}^{*}}\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}(\hat{\pi}_{i}-\tilde{\pi}_{i}).$ (19)
We first consider the third term. From Lemma 9 and using mean value theorem and that $\phi^{\prime}(\cdot)$ is uniformly bounded, we get with probability converging to $1$, for all $i$,
$|\hat{\pi}_{i}-\pi_{i}^{*}|\lesssim\left|X_{i}^{\top}(\hat{\gamma}-\gamma^{*})\right|.$
Treating $|\tilde{\pi}_{i}-\pi_{i}^{*}|$ analogously, and using that $\frac{T_{i}}{\pi_{i}^{*}},\frac{T_{i}}{\tilde{\pi}_{i}},\frac{T_{i}}{\hat{\pi}_{i}},\hat{\mu}_{i}$ are, with probability converging to $1$, all bounded, we can bound the third term via the following chain of inequalities:
$\frac{1}{n}$ $\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-\frac{T_{i}}{\pi_{i}^{*}}\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}(\hat{\pi}_{i}-\tilde{\pi}_{i})\lesssim\frac{1}{n}\sum_{i=1}^{n}\left|X_{i}^{\top}(\hat{\gamma}-\gamma^{*})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})\right|$
$\leq\frac{1}{n}\left\|\mathbf{X}(\hat{\gamma}-\gamma^{*})\right\|_{2}^{2}+\frac{1}{n}\left\|\mathbf{X}(\hat{\gamma}-\gamma^{*})\right\|_{2}\left\|\tilde{\mathbf{X}}(\tilde{\gamma}-\hat{\gamma})\right\|_{2},$
where for the last inequality we decompose $\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})=X_{i}^{\top}(\hat{\gamma}-\gamma^{*})+\tilde{X}_{i}^{\top}(\tilde{\gamma}-\hat{\gamma})$ and employ Cauchy-Schwarz inequality. Then in light of Lemma 4, Condition 8 and $s_{\pi}=o(\sqrt{n}/\sqrt{\log p\log p\vee n})$, we can show the third term to be of order $o_{\mathbb{P}}(1)$.
We now focus on the first term. Using Taylor’s theorem and Lemma 9, we have that for some $\iota_{i},\iota_{i}^{\prime}\in[0,1]$ for $i=1,\ldots,n$,
$\mathrm{I}=$ $\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}^{\top}(\hat{\gamma}-\tilde{\gamma})$
$+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime\prime}(\iota_{i}X_{i}^{\top}\hat{\gamma}+(1-\iota_{i})\tilde{X}_{i}^{\top}\tilde{\gamma})(\tilde{X}_{i}^{\top}(\hat{\gamma}-\tilde{\gamma}))^{2}+o_{\mathbb{P}}(1)$
and
$\mathrm{II}=$ $\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}^{\top}(\hat{\gamma}-\tilde{\gamma})$
$+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime\prime}(\iota_{i}^{\prime}X_{i}^{\top}\hat{\gamma}+(1-\iota_{i}^{\prime})\tilde{X}_{i}^{\top}\tilde{\gamma})(\tilde{X}_{i}^{\top}(\hat{\gamma}-\tilde{\gamma}))^{2}+o_{\mathbb{P}}(1).$
Applying an analogous analysis as above, we can control the second terms in the above two decompositions to be of order $o_{\mathbb{P}}(1/\sqrt{n})$. Thus, to prove the desired result, it remains to show that the first terms in the above two decompositions are of order $o_{\mathbb{P}}(1/\sqrt{n})$ as well. Applying Hölder’s inequality, this is a direct consequence of Lemma 4 and the following lemma.
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}$
@@
By conditioning on $(\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}})$, $\{\frac{T_{i}}{\pi_{i}^{*}}-1,i=1,\cdots,n\}$ of the main dataset $\mathcal{D}$ are independent mean-zero and bounded random variables. Then as a direct consequence of Hoeffding’s inequality, we obtain the first inequality. For the second inequality, following the same proof as in Wang and Shah [2020, Lemma 14], we get, with probability converging to $1$,
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime}(X_{i}^{\top}\hat{\gamma})X_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p}{n}}.$
@@
Finally, using that the augmented entries of $\tilde{X}_{i}$ are i.i.d. random variables drawn uniformly from $[-1,1]$ and are independent of $(\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}})$, the second inequality is again a direct consequence of Hoeffding’s inequality. ∎
A.6 Proof of Theorem 6
A.6.1 Proof of Theorem 6(i)
Using the same analysis as in Lemma 1, we have that with probability converging to $1$, the $\tilde{\gamma}$ such that $\pi_{i}^{*}\equiv\phi(\tilde{X}_{i}^{\top}\tilde{\gamma})$ is a feasible solution to the convex program. Now we have the decomposition
@@ A.6.1 Proof of Theorem 6(i)
$\hat{\tau}_{\textrm{DCal}}-\tau^{*}=$ ($\ \underset{\eqqcolon\mathrm{I}}{\underbrace{(\hat{\tau}-\tau^{*})\left(1-\frac{\sum_{i=1}^{n}T_{i}(T_{i}-\tilde{\pi}_{i})}{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}_{i})^{2}}\right)}}+\underset{\eqqcolon\mathrm{II}}{\underbrace{\frac{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}(X_{i}))(r(X_{i})-\hat{r}(X_{i}))}{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}(X_{i}))^{2}}}}$

### v4

@@ A.5 Proof of Theorem 5
We begin with the following lemma on the negligible effect of trimming in $\hat{\pi}$ when $s_{\pi}=o(\sqrt{n}/\log p)$:
@@
Under the conditions of Theorem 5, we have that by choosing $M_{\gamma}\asymp 1$ sufficiently large, with probability converging to $1$, for all $i$,
$\hat{\pi}_{i}\equiv\phi(X_{i}^{\top}\hat{\gamma})\quad\&\quad\hat{\pi}_{\mathrm{aux},i}\equiv\phi(X_{\mathrm{aux},i}^{\top}\hat{\gamma}).$
@@
where for the last inequality we use Condition 7 and that $s_{\pi}=o(\sqrt{n}/\log p)$. Thus the statement of this lemma is proved. ∎
Armed with the above lemma, we can see that in the large sample limit, $\hat{\tau}_{\textrm{SCal},r}$ is numerically equivalent to the estimator by redefining $\hat{\pi}\coloneqq\phi(X^{\top}\hat{\gamma})$, which allows us to apply the same analysis in Wang and Shah [2024] to understand the asymptotic property of $\hat{\bm{\mu}}$. Specifically, we have
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\gamma^{*})\tilde{X}_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}$
@@
By conditioning on $(\tilde{\mathbf{X}},\hat{\bm{\mu}})$, $\{\frac{T_{i}}{\pi_{i}^{*}}-1,i=1,\cdots,n\}$ of the main dataset $\mathcal{D}$ are independent mean-zero and bounded random variables. Then as a direct consequence of Hoeffding’s inequality, we obtain that for some constants $c_{1},c_{2}>0$, almost surely
@@
$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\gamma^{*})\tilde{X}_{i}\right\|_{\infty}\leq c_{1}\max_{i}\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\sqrt{\frac{\log p\vee n}{n}}\;\bigg{|}\;\tilde{\mathbf{X}},\hat{\bm{\mu}}\right)\leq c_{2}(p\vee n)^{-1}.$
Using above, the first inequality follows from Lemma 2.
For the second inequality, from Lemmas 8 and 9, we get, with probability converging to $1$,
$\left\|\frac{1}{n_{\mathrm{aux}}}\sum_{i=1}^{n_{\mathrm{aux}}}\frac{T_{i,\mathrm{aux}}(Y_{i,\mathrm{aux}}-\hat{r}_{i,\mathrm{aux}})}{\hat{\pi}_{i,\mathrm{aux}}^{2}}\phi^{\prime}(X_{i,\mathrm{aux}}^{\top}\hat{\gamma})X_{i,\mathrm{aux}}-\frac{1}{n}\sum_{i=1}^{n}\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\hat{\gamma})X_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p}{n}}.$
It also follows from the proof of Lemma 8 that with probability converging to $1$,
$\left\|\frac{1}{n_{\mathrm{aux}}}\sum_{i=1}^{n_{\mathrm{aux}}}\frac{T_{i,\mathrm{aux}}(Y_{i,\mathrm{aux}}-\hat{r}_{i,\mathrm{aux}})}{\hat{\pi}_{i,\mathrm{aux}}^{2}}\phi^{\prime}(X_{i,\mathrm{aux}}^{\top}\hat{\gamma})X_{i,\mathrm{aux}}-\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}\phi^{\prime}(X_{i}^{\top}\hat{\gamma})X_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p}{n}}.$
@@
In light of the above inequalities, and using that the augmented entries of $\tilde{X}_{i}$ are i.i.d. random variables drawn uniformly from $[-1,1]$ and are independent of $(\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}})$, the second inequality is again a direct consequence of Hoeffding’s inequality. ∎
Proof of Theorem 5.
Without loss of generality, in this proof we assume throughout that $p\geq n$, so that $\mathbf{X}$ and $\tilde{\mathbf{X}}$ are equivalent. The proof of the case $p<n$ can be derived via an analogous argument. To prove the desired statement, it boils down to proving the following term to be $o_{\mathbb{P}}(1)$.
$Q\coloneqq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\pi}_{i}}-\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\pi_{i}^{*}}\right).$ (29)
We apply the following decomposition:
$Q$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\pi_{i}^{*}\tilde{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}\right)(\pi_{i}^{*}-\tilde{\pi}_{i})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}(\pi_{i}^{*}-\tilde{\pi}_{i})$
@@
$=\underset{\mathrm{I}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\pi_{i}^{*}\tilde{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})}}+\underset{\mathrm{II}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\pi_{i}^{*}\tilde{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}\right)(\pi_{i}^{*}-\hat{\pi}_{i})}}$ (30)
$\quad-\underset{\mathrm{III}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}(\pi_{i}^{*}-\tilde{\pi}_{i})}}.$
We first consider $\mathrm{I}$, which we further decompose as
$\mathrm{I}$ $=\underset{\mathrm{I}_{1}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})}}+\underset{\mathrm{I}_{2}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\pi_{i}^{*}\tilde{\pi}_{i}}-\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})}}$
$\quad-\underset{\mathrm{I}_{3}}{\underbrace{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})}}$
Using Taylor expansion and Lemma 9, we have that for some $\iota_{i}\in[0,1]$ for $i=1,\ldots,n$,
$\mathrm{I}_{1}=$ $\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime}(X_{i}^{\top}\hat{\gamma})X_{i}^{\top}(\hat{\gamma}-\tilde{\gamma})$
$+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime\prime}(\iota_{i}X_{i}^{\top}\hat{\gamma}+(1-\iota_{i})X_{i}^{\top}\tilde{\gamma})(X_{i}^{\top}(\hat{\gamma}-\tilde{\gamma}))^{2}.$
For the second term, using that with probability converging to $1$, $\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\hat{\pi}_{i}^{2}},\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}$ are all bounded below by some constant and that $\phi^{\prime\prime}$ is uniformly bounded, as well as Lemma 4 and $s_{\pi}=o(\sqrt{n}/\sqrt{\log p\log p\vee n})$, we have the second term in the above decomposition is of order $o_{\mathbb{P}}(1)$. For the first term, applying Hölder’s inequality, it is a direct consequence of Lemmas 4 and 10.
For $\mathrm{I}_{2}$ and $\mathrm{I}_{3}$, from Lemma 9, using mean value theorem and that $\phi^{\prime}(\cdot)$ is uniformly bounded, we get with probability converging to $1$, for all $i$,
$|\hat{\pi}_{i}-\pi_{i}^{*}|\lesssim\left|X_{i}^{\top}(\hat{\gamma}-\gamma^{*})\right|\quad\&\quad|\hat{\pi}_{i}-\tilde{\pi}_{i}|\lesssim\left|X_{i}^{\top}(\hat{\gamma}-\tilde{\gamma})\right|.$
Moreover, using that $\hat{\gamma}$ and the main data set $\mathcal{D}$ are independent, Condition 7 and a Bernstein’s inequality (see e.g. (23)), we can prove that with probability converging to $1$,
$\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}(\hat{\gamma}-\gamma^{*}))^{2}=O\left(s_{\pi}\frac{\log p}{n}\right).$
In light of the above two inequalities, the control of $\mathrm{I}_{2}$ and $\mathrm{I}_{3}$ follows from the same analysis as the second term in the decomposition of $\mathrm{I}_{1}$. Putting together we prove that $\mathrm{I}=o_{\mathbb{P}}(1)$. Using an analogous argument we also prove that $\mathrm{I}\mathrm{I}=o_{\mathbb{P}}(1)$.
It now remains to consider $\mathrm{I}\mathrm{I}\mathrm{I}$. Using Taylor expansion, we have that for some $\iota_{i}^{\prime}\in[0,1]$ for $i=1,\ldots,n$,
$\mathrm{I}\mathrm{I}\mathrm{I}=$ $\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\tilde{\pi}_{i}}\phi^{\prime}(X_{i}^{\top}\gamma^{*})X_{i}^{\top}(\gamma^{*}-\tilde{\gamma})$
$+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi_{i}^{*}}-1\right)\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\phi^{\prime\prime}(\iota_{i}^{\prime}X_{i}^{\top}\gamma^{*}+(1-\iota_{i}^{\prime})X_{i}^{\top}\tilde{\gamma})(X_{i}^{\top}(\gamma^{*}-\tilde{\gamma}))^{2}.$
Applying an analogous analysis as above, we can show that the second terms in the above two displays are of order $o_{\mathbb{P}}(1)$. Thus, to prove the desired result, it remains to show that the first terms in the above two displays are also of order $o_{\mathbb{P}}(1)$. Applying Hölder’s inequality, this is a direct consequence of Lemmas 4 and 10. ∎
Appendix B Proofs of the results concerning Section 4
B.1 Proof of Theorem 6
B.1.1 Proof of Theorem 6 (i)
Using the same analysis as in Lemma 1, there exists some $\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{\pi}\asymp 1$ such that with probability converging to $1$, the $\tilde{\gamma}^{*}$ such that $\pi_{i}^{*}\equiv\phi(\tilde{X}_{i}^{\top}\tilde{\gamma}^{*})$ for all $i$ is a feasible solution to (14). We now apply the decomposition
$\hat{\tau}_{\mathrm{DCal}}-\tau^{*}=$ ($\ \underset{\eqqcolon\,\mathrm{I}}{\underbrace{(\hat{\tau}-\tau^{*})\left(1-\frac{\sum_{i=1}^{n}T_{i}(T_{i}-\tilde{\pi}_{i})}{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}_{i})^{2}}\right)}}+\underset{\eqqcolon\,\mathrm{II}}{\underbrace{\frac{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}(X_{i}))(r(X_{i})-\hat{r}(X_{i}))}{\sum_{i=1}^{n}(T_{i}-\tilde{\pi}(X_{i}))^{2}}}}$